OpenStudy (anonymous):

-7(x-4)+8x=12+14 would x=11?

OpenStudy (anonymous):

no it wouldn't because if you use a calculator to do this it equals to 39 and 12 + 14 is 26.

OpenStudy (anonymous):

I'm sorry I still don't get it, what is x equal to, I've been stuck on this for a while now :/

OpenStudy (anonymous):

okay so you have to plug in 11 first so it would be -7(11-4)+8(11)= 39 and 12+14=26. 39 and 26 arnt equal so that equation is wrong

OpenStudy (anonymous):

to make this easier copy and paste this -7(11-4)+8(11)= to google and the answer will show up which is 39 so that half is 39 and the 12+ 14 = that equals 26 those arnt equal so its wrong. I hoped this helped :)

OpenStudy (anonymous):

I don't get it, math is a very hard subject for me....

OpenStudy (anonymous):

so x= -2

OpenStudy (anonymous):